When you drop a matchbox onto the floor, there is a 15% probability of it landing on the side. Suppose the matchbox is dropped onto the floor 8 times. 
a) What is the probability of it NOT landing on the side on any the 8 times it's dropped? 
b) What is the probability that it will land on the side some of the times it's dropped? (one or more of the 8 times)
Answer #1

P(Match box lands on a side) = 0.15
a)

P(Match box does not land on it's side) = 1 - 0.15 = 0.85
P(Match box does not land on it's side all 8 times) = 0.858 = 0.2725

b)
P(Match box lands on its side atleast once) = 1 - P(Match box does not land on its side all 8 times) = 1 - 0.2725 = 0.7275